OpenStudy (anonymous):

2(2k - 1) = 4(k-2) SHOW WORK PLEASE

OpenStudy (anonymous):

distribute the 2 and 4 first

OpenStudy (anonymous):

2(2k-1)=4(k-2) distribute: 4k-2=4k-8 bring numbers to one side: 4k+6=4k answer is no solution